# Integrable deformations of principal chiral model from

### v1

@@ 2.4 $U$-matrices and integrability
Consider solutions to the moment map equation $\mu_{0}=0$ (2.35) with $\mu_{0}$ (2.31):
$\begin{array}[]{c}\displaystyle{\partial_{\bar{z}}\Phi-{k}\partial_{x}\bar{A}+[\bar{A},\Phi]=\sum_{a=1}^{n}S^{a}\delta(z-z_{a},\bar{z}-\bar{z}_{a})}\end{array}$ (2.36)
Here we study the case $\Sigma$ is an elliptic curve and $G={\rm GL}(N,{\mathbb{C}})$, so that $\Phi$, $\bar{A}$ and the set of $S^{a}$, $a=1,...,n$ become matrices from ${\rm Mat}(N,\mathbb{C})$. Notice that using the gauge group transformations one can make ${\bar{A}}$ be diagonal (in the general case ${\bar{A}}$ is transformed to an element from the Cartan subalgebra of ${\mathfrak{g}}$). Perform a conjugation of both sides with $\exp((z-{\bar{z}}){\bar{A}})$. Then we get (2.36) in the ”holomorphic gauge”:
@@
where $\tilde{\Phi}=e^{-(z-{\bar{z}}){\bar{A}}}\Phi e^{(z-{\bar{z}}){\bar{A}}}$. Let ${k}$ be a constant for simplicity. More complicated cases are considered below in the next subsection. Solution of this equation requires some boundary conditions and a gauge fixation of ${\bar{A}}$. The gauge fixation of ${\bar{A}}$ depends on topological properties of the Higgs bundle. In fact, the Higgs field $\tilde{\Phi}$ is a section of the de3fined above bundle ${\rm End}(E^{aff}))$. The mentioned above topological property is the degree of the underlying bundle $E$. When ${\rm deg}(E)=0$ the matrix ${\bar{A}}$ is transformed to the constant diagonal matrix. The diagonal elements become dynamical variables. Then the solution of (2.37) yields the $U$-matrix of the field analogue of the Calogero-Moser model or its spin and multispin (multipole) generalizations [16, 34].
@@
Explicit expression for $U$-matrix in elliptic functions will be given in the next Section, see (3.25). Here we emphasize that the equation (2.38) or (2.39) is independent of derivatives with respect to loop variable. That is (2.38) has the same form as in the finite-dimensional case. Therefore, its solution has the same form as in mechanics but the residues $S^{a}$ are now considered as functions of $x$.
@@
Moreover, in our case the classical $r$-matrix is skew-symmetric (${\bf r}_{21}(w,z|x)=-{\bf r}_{12}(z,w|x)$), depends on the difference of spectral parameters ${\bf r}_{12}(z,w|x)={\bf r}_{12}(z-w|x)$ and it is non-dynamical: ${\bf r}_{12}(z-w|x)=r_{12}(z-w)$. The latter $r$-matrix satisfies the standard classical Yang-Baxter equation for non-dynamical skew-symmetric $r$-matrices
@@
The $U$-matrix satisfies the classical $r$-matrix structure (2.41).
This Proposition guarantees integrability in the following sense. Existence of the classical $r$-matrix structure (2.41) is a sufficient condition for the Poisson commutativity $\{{\rm tr}(T^{k}(z,2\pi)),{\rm tr}(T^{m}(w,2\pi))\}=0$ of the traces of powers of the matrices
@@ 2.5 Cocentral charge as twist function
In the above analysis of $U$-matrices we assumed $k$ be a constant. At the same time it satisfies the equation 2 from (2.35)
@@
$\begin{array}[]{c}\displaystyle{\mathop{\hbox{Res}}\limits\limits_{z=w_{b}}k(z)=s_{b}\,,\quad b=1,...,m}\end{array}$ (2.46)
@@
$\begin{array}[]{c}\displaystyle{{k}(z)=\prod_{b=1}^{m}\frac{z-y_{b}}{z-w_{b}}}\end{array}$ (2.47)
@@
This solution satisfies (2.34).
2. $\Sigma$ is an elliptic curve. In this case solutions of the moment equation (2.33) should be double-periodic. We take
@@ 3.1 $R$-matrices
It can be shown (see e.g. [17]) that any solution of (3.1) obeying the properties (3.4) and (3.5) satisfies also the quantum Yang-Baxter equation (1.9). This is why we indeed deal with quantum $R$-matrices. Let us briefly describe families of $R$-matrices which we are dealing with. Before considering non-trivial examples let us write down the most simple one, which is the (properly normalized) rational Yang’s $R$-matrix:
@@
and the basis matrices are $T_{00}=1_{2}=\sigma_{0}$, $T_{10}=-\sigma_{3}$, $T_{10}=\sigma_{1}$ and $T_{11}=\sigma_{2}$, where $\{\sigma_{a}\}$, $a=0,1,2,3,4$ are the Pauli matrices:
@@
$\begin{array}[]{c}\displaystyle{R_{12}^{\hbar}(z)=\frac{1}{2}\Big{(}\varphi_{00}\,\sigma_{0}\otimes\sigma_{0}+\varphi_{01}\,\sigma_{1}\otimes\sigma_{1}+\varphi_{11}\,\sigma_{2}\otimes\sigma_{2}+\varphi_{10}\,\sigma_{3}\otimes\sigma_{3}\Big{)}\,,}\end{array}$ (3.21)
$\begin{array}[]{c}\displaystyle{\varphi_{00}=\phi(z,\frac{\hbar}{2})\,,\quad\varphi_{10}=\phi(z,\frac{1}{2}+\frac{\hbar}{2})\,,\quad\varphi_{01}=e^{\pi\imath z}\phi(z,\frac{\tau}{2}+\frac{\hbar}{2})\,,\quad\varphi_{11}=e^{\pi\imath z}\phi(z,\frac{1+\tau}{2}+\frac{\hbar}{2})\,.}\end{array}$ (3.22)
In $4\times 4$ form it is as follows:
@@
$\begin{array}[]{c}R^{\hbar}(z)=\left(\begin{array}[]{cccc}\coth(z)+\coth(\hbar)&0&0&0\vphantom{\Big{|}}\\ 0&\sinh^{-1}(\hbar)&\sinh^{-1}(z)&0\vphantom{\Big{|}}\\ 0&\sinh^{-1}(z)&\sinh^{-1}(\hbar)&0\vphantom{\Big{|}}\\ -4\,e^{-2\Lambda}\sinh(z+\hbar)&0&0&\coth(z)+\coth(\hbar)\vphantom{\Big{|}}\end{array}\right)\,.\end{array}$ (3.26)
@@ 3.3 Euler-Arnold tops, Gaudin models and Painlevé-Schlesinger systems
$\begin{split}&L(S^{a},z-z_{a})L(S^{b},z-z_{b})=L(S^{a}I^{ab}(S^{b}),z-z_{a})+L(I^{ba}(S^{a})S^{b},z-z_{b})+\\ &+\frac{{\rm tr}(S^{b})}{N}M(S^{a},z-z_{a})+\frac{{\rm tr}(S^{a})}{N}M(S^{b},z-z_{b})+\frac{1}{N}{\rm tr}_{23}({m_{23}(z_{a}-z_{b})S^{a}_{2}S^{b}_{3}})\end{split}$ (3.57)
@@ 4.1 Higher rank Landau-Lifshitz equation
Let us briefly recall the construction of higher rank Landau-Lifshitz model from [^7] There is a small difference between [2] and the below given construction since in this paper we use the Zakharov-Shabat equation in the form (1.5), while in [2] it is $\partial_{t}U(z)-k\partial_{x}V(z)=[U(z),V(z)]$, and $k$ is fixed as $k=1$. [2]. As was argued previously, $U$-matrix has the same form as in the finite-dimensional mechanics (3.39):
@@
where $c\in{\mathbb{C}}$ is a constant and $T\in{\rm Mat}(N,\mathbb{C})$ – some dynamical matrix-valued variable. It will be fixed below. The constant $c$ appears in the additional property, which we require for the matrix $S$:
@@ 4.3.2 Equations of motion for the first flows
$\begin{array}[]{l}\displaystyle{\partial_{t_{a}}U(z)-k\partial_{x}V^{a}(z)=[U(z),V^{a}(z)]}\end{array}$ (4.33)
@@
This one system is a generalization of the equations (1.4). For the rational XXX $r$-matrix $r_{12}(z)=P_{12}/z$ (1.4) is reproduced from (4.38).
@@ 4.3.3 Equations of motion for the second flows
$\begin{array}[]{c}\partial_{t_{a}}S^{a}=\\ \\ =k\partial_{x}(E(S^{a})S^{a}+S^{a}E(S^{a}))+kc\partial_{x}T^{a}+c\left[S^{a},E(T^{a})\right]+c\left[E(S^{a}),T^{a}\right]+c\left[I^{ab}(S^{b}),T^{a}\right]-\\ \\ -\frac{2{\rm tr}(S^{a})}{N}\left[S^{a},J(S^{a})\right]+S^{a}E(\left[I^{ab}(S^{b}),S^{a}\right])+E(\left[I^{ab}(S^{b}),S^{a}\right])S^{a}+E(S^{a})\left[I^{ab}(S^{b}),S^{a}\right]+\\ \\ +\left[I^{ab}(S^{b}),S^{a}\right]E(S^{a})+S^{a}I^{ab}(\left[S^{b},I^{ab}(S^{a})\right])+I^{ab}(\left[S^{b},I^{ab}(S^{a})\right])S^{a}\,,\end{array}$ (4.42)
@@
$\begin{array}[]{c}\displaystyle{V^{a}(z)=c\frac{S^{a}}{(z-z_{a})^{2}}-\frac{c^{-1}k}{z-z_{a}}\left[S^{a},\partial_{x}S^{a}\right]+\frac{cS^{b}}{(z-z_{a})\cdot(z_{a}-z_{b})}\,.}\end{array}$ (4.46)
@@ 4.3.4 Minimal orbits
$\begin{array}[]{c}\displaystyle{S^{a}=\xi^{a}\otimes\eta^{a}\,,\quad S^{b}=\xi^{b}\otimes\eta^{b}}\end{array}$ (4.48)
@@
The upper given identities allows to simplify the equation (4.43). Namely, we get (here we have already substituted $T^{a}$ from (4.45)):

### v4

@@ 2.1 $U$-matrices and integrability
Now, let us proceed to the field case. The moment map equation takes the form [^5] In the Appendix (2.2) appears as $\mu_{0}=0$ in (A.35) with $\mu_{0}$ (A.31).:
$\begin{array}[]{c}\displaystyle{\partial_{\bar{z}}\Phi-{k}\partial_{x}\bar{A}+[\bar{A},\Phi]=\sum_{a=1}^{n}S^{a}\delta(z-z_{a},\bar{z}-\bar{z}_{a})\,.}\end{array}$ (2.2)
Equation (2.2) is the field generalization of the moment map equation (2.1) Here we study the case $\Sigma$ is an elliptic curve and $G={\rm GL}(N,{\mathbb{C}})$, so that $\Phi$, $\bar{A}$ and the set of $S^{a}$, $a=1,...,n$ become matrices from ${\rm Mat}(N,\mathbb{C})$. Notice that using the gauge group transformations one can make ${\bar{A}}$ be diagonal (in the general case ${\bar{A}}$ is transformed to an element from the Cartan subalgebra of ${\mathfrak{g}}$). Perform a conjugation of both sides with $\exp((z-{\bar{z}}){\bar{A}})$. Then we get (2.2) in the ”holomorphic gauge”:
@@
where $\tilde{\Phi}=e^{-(z-{\bar{z}}){\bar{A}}}\Phi e^{(z-{\bar{z}}){\bar{A}}}$. Let ${k}$ be a constant for simplicity. More complicated cases are considered below in the next subsection. Solution of this equation requires some boundary conditions and a gauge fixation of ${\bar{A}}$. The gauge fixation of ${\bar{A}}$ depends on topological properties of the Higgs bundle. In fact, the Higgs field $\tilde{\Phi}$ is a section of the affine Higgs bundle ${\rm End}(E^{aff})$, see Appendix. The mentioned above topological property is the degree of the underlying bundle $E$. When ${\rm deg}(E)=0$ the matrix ${\bar{A}}$ can be transformed to the constant diagonal matrix. The diagonal elements become dynamical variables. Then the solution of (2.3) yields the $U$-matrix of the field analogue of the Calogero-Moser model or its spin and multispin (multipole) generalizations [17, 35].
@@
Explicit expression for $U$-matrix in elliptic functions will be given in the next Section, see (3.25). Here we emphasize that the equation (2.4) or (2.5) is free of derivatives with respect to the loop variable $x$. That is (2.4) has the same form as in the finite-dimensional case. Therefore, its solution has the same form as in mechanics but the residues $S^{a}$ are now considered as functions of $x$.
@@
Moreover, in our case the classical $r$-matrix is skew-symmetric (${\bf r}_{21}(w,z|x)=-{\bf r}_{12}(z,w|x)$), depends on the difference of spectral parameters ${\bf r}_{12}(z,w|x)={\bf r}_{12}(z-w|x)$ and it is non-dynamical: ${\bf r}_{12}(z-w|x)=r_{12}(z-w)$. This type $r$-matrix satisfies the standard classical Yang-Baxter equation for non-dynamical skew-symmetric $r$-matrices
@@
The matrix $U(z)$ (2.9) satisfies the classical $r$-matrix structure (2.7).
This Proposition guarantees integrability in the following sense. Existence of the classical $r$-matrix structure (2.7) is a sufficient condition for the Poisson commutativity $\{{\rm tr}(T^{k}(z,2\pi)),{\rm tr}(T^{m}(w,2\pi))\}=0$ of the traces of powers of the monodromy matrices
@@ 2.2 Cocentral charge as twist function
In the above analysis of $U$-matrices we assumed $k$ be a constant. At the same time it satisfies the equation 2 from (A.35)
@@
$\begin{array}[]{c}\displaystyle{\mathop{\hbox{Res}}\limits\limits_{z=w_{b}}k(z)=s_{b}\,,\quad b=1,...,m,}\end{array}$ (2.12)
@@
$\begin{array}[]{c}\displaystyle{{k}(z)=\prod_{b=1}^{m}\frac{z-y_{b}}{z-w_{b}},}\end{array}$ (2.13)
@@
This solution satisfies (A.34).
2. $\Sigma$ is an elliptic curve. In this case solutions of the moment equation (A.33) should be double-periodic. We take
@@ 3.1 $R$-matrices
It can be shown (see e.g. [18]) that any solution of (3.1) obeying the properties (3.4) and (3.5) satisfies also the quantum Yang-Baxter equation (1.9). This is why we indeed deal with quantum $R$-matrices. More precisely, we deal with a subset of solutions of the quantum Yang-Baxter equation (1.9), which also satisfy the AYBE (3.1). Let us briefly describe families of $R$-matrices which satisfy all required properties. Before considering non-trivial examples let us write down the most simple one, which is the (properly normalized) rational Yang’s $R$-matrix:
@@
and the basis matrices are $T_{00}=1_{2}=\sigma_{0}$, $T_{10}=-\sigma_{3}$, $T_{01}=\sigma_{1}$ and $T_{11}=\sigma_{2}$, where $\{\sigma_{a}\}$, $a=0,1,2,3,4$ are the Pauli matrices:
@@
$\begin{array}[]{c}\displaystyle{R_{12}^{\hbar}(z)=\varphi_{00}\,\sigma_{0}\otimes\sigma_{0}+\varphi_{01}\,\sigma_{1}\otimes\sigma_{1}+\varphi_{11}\,\sigma_{2}\otimes\sigma_{2}+\varphi_{10}\,\sigma_{3}\otimes\sigma_{3}\,,}\end{array}$ (3.21)
$\begin{array}[]{c}\displaystyle{\varphi_{00}=\phi(z,\hbar)\,,\quad\varphi_{10}=\phi(z,\hbar)\,,\quad\varphi_{01}=e^{\pi\imath z}\phi(z,\hbar)\,,\quad\varphi_{11}=e^{\pi\imath z}\phi(z,\hbar)\,.}\end{array}$ (3.22)
In the form of $4\times 4$ matrix it is as follows:
@@
$\begin{array}[]{c}R^{\hbar}(z)=\left(\begin{array}[]{cccc}\coth(z)+\coth(\hbar)&0&0&0\vphantom{\Big{|}}\\ 0&\sinh^{-1}(\hbar)&\sinh^{-1}(z)&0\vphantom{\Big{|}}\\ 0&\sinh^{-1}(z)&\sinh^{-1}(\hbar)&0\vphantom{\Big{|}}\\ -4\,e^{-2\Lambda}\sinh(z+\hbar)&0&0&\coth(z)+\coth(\hbar)\vphantom{\Big{|}}\end{array}\right)\,,\end{array}$ (3.26)
@@ 3.3 Euler-Arnold tops, Gaudin models and Painlevé-Schlesinger systems
$\begin{split}&L(S^{a},z-z_{a})L(S^{b},z-z_{b})=L(S^{a}I^{ab}(S^{b}),z-z_{a})+L(I^{ba}(S^{a})S^{b},z-z_{b})+\\ &+\frac{{\rm tr}(S^{b})}{N}M(S^{a},z-z_{a})+\frac{{\rm tr}(S^{a})}{N}M(S^{b},z-z_{b})+\frac{1}{N}{\rm tr}_{23}({m_{23}(z_{a}-z_{b})S^{a}_{2}S^{b}_{3}})\,.\end{split}$ (3.57)
@@ 4.1 Higher rank Landau-Lifshitz equation
Let us briefly recall the construction of higher rank Landau-Lifshitz model from [^8] There is a small difference between [2] and the below given construction since in this paper we use the Zakharov-Shabat equation in the form (1.5), while in [2] it is $\partial_{t}U(z)-k\partial_{x}V(z)=[U(z),V(z)]$, and $k$ is fixed as $k=1$. [2]. As was explained previously, $U$-matrix has the same form as in the finite-dimensional mechanics (3.39):
@@
where $c\in{\mathbb{C}}$ is a constant and $T\in{\rm Mat}(N,\mathbb{C})$ – some dynamical matrix-valued variable. It will be defined below. The constant $c$ appears in the additional property, which we require for the matrix $S$:
@@ 4.3.2 Equations of motion for the first flows
$\begin{array}[]{l}\displaystyle{\partial_{t_{a}}U(z)-k\partial_{x}V^{a}(z)=[U(z),V^{a}(z)]\,,}\end{array}$ (4.33)
@@
This one system is a generalization of the equations (1.4). For the rational XXX $r$-matrix $r_{12}(z)=P_{12}/z$ the equations (1.4) are reproduced from (4.38).
@@ 4.3.3 Equations of motion for the second flows
$\begin{array}[]{c}\partial_{t_{a}}S^{a}=\\ \\ =k\partial_{x}(E(S^{a})S^{a}+S^{a}E(S^{a}))+kc\partial_{x}T^{a}+c\left[S^{a},E(T^{a})\right]+c\left[E(S^{a}),T^{a}\right]+c\left[I^{ab}(S^{b}),T^{a}\right]-\\ \\ -\frac{2{\rm tr}(S^{a})}{N}\left[S^{a},J(S^{a})\right]+S^{a}E(\left[I^{ab}(S^{b}),S^{a}\right])+E(\left[I^{ab}(S^{b}),S^{a}\right])S^{a}+E(S^{a})\left[I^{ab}(S^{b}),S^{a}\right]+\\ \\ +\left[I^{ab}(S^{b}),S^{a}\right]E(S^{a})+S^{a}I^{ab}(\left[S^{b},I^{ab}(S^{a})\right])+I^{ab}(\left[S^{b},I^{ab}(S^{a})\right])S^{a}\end{array}$ (4.42)
@@
$\begin{array}[]{c}\displaystyle{V^{a}(z)=c\frac{S^{a}}{(z-z_{a})^{2}}-\frac{c^{-1}k}{z-z_{a}}\left[S^{a},\partial_{x}S^{a}\right]+\frac{cS^{b}}{(z-z_{a})(z_{a}-z_{b})}\,.}\end{array}$ (4.46)
@@ 4.3.4 Minimal orbits
$\begin{array}[]{c}\displaystyle{S^{a}=\xi^{a}\otimes\eta^{a}\,,\quad S^{b}=\xi^{b}\otimes\eta^{b}\,,}\end{array}$ (4.48)
@@
The upper identities allows to simplify the equation (4.43). Namely, we get (here we have already substituted $T^{a}$ from (4.45)):
@@
Let us mention that the Hamiltonian structure for the derived equations is unknown. It is an interesting ad important problem. In fact, even for a single marked point case (i.e. for the Landau-Lifshitz equation (4.11)) it is known for the special coadjoint orbit (4.12) only. In this case (4.11) reduces to (4.20).
@@ 4.5 Deformation by the twist function
5 Appendix: the Higgs bundles and Hitchin systems
Here we briefly describe main steps of construction of integrable systems through the Hitchin approach. Main idea is to perform the Hamiltonian reduction starting from the Higgs field defined on the corresponding Higgs bundle. The moment map equation is a constrain which projects the Higgs field to the Lax matrix. Similarly, in the case of the field theory one should change the Higgs bundle to the affine Higgs bundle. Then we deal with the Lax connection (the $U$-matrix), and it satisfies a field analogue of the moment map equation.
5.1 Preliminaries on loop groups and loop algebras
$\langle X\,,\,Y\rangle=\oint(X,Y)\frac{dy}{y}\,.$
Consider its central extension
$\hat{L}({\mathfrak{g}})=\{(X(x),\kappa)\}\,,~{}~{}\kappa\in{\mathbb{C}}\,.$ (A.1)
The commutator of $\hat{L}({\mathfrak{g}})$ assumes the form
where $[(X_{1},X_{2})]_{0}$ is a commutator on ${\mathfrak{g}}$, The dual to $\hat{L}({\mathfrak{g}})$ is the Lie coalgebra
$\hat{L}^{*}({\mathfrak{g}})=\{\mathcal{Y}=(Y,{k})\sim({k}\partial+Y)\}\,.$ (A.2)
It is defined through the pairing
$(\mathcal{X},\mathcal{Y})=\langle X,Y\rangle+\kappa{k}\,,$ (A.3)
where ${k}$ is the cocenter.
$L(G)=G\otimes{\mathbb{C}}(y))=\left\{\,\sum_{m}g_{m}y^{m}\,,~{}g_{m}\in G\,\right\}\,.$ (A.4)
$(g,\zeta)\times(g^{\prime},\zeta^{\prime})=\left(gg^{\prime},\zeta\zeta^{\prime}{\mathcal{C}}(g,g^{\prime})\right),$ (A.5)
The adjoint action of $f\in L(G)$ on $\hat{L}({\mathfrak{g}})$ is defined as
${\rm Ad}_{f}\mathcal{X}={\rm Ad}_{f}(X,\kappa)=(fXf^{-1}\,,\,\kappa+\langle f^{-1}\partial f,X\rangle)\,.$ (A.6)
Then the coadjoint action of $L(G)$ assumes the form
${\rm Ad}^{*}_{f}\mathcal{Y}={\rm Ad}^{*}_{f}(Y,{k})=(f^{-1}Yf+{k}f^{-1}\partial f\,,\,{k})\,.$ (A.7)
5.2 Briefly on Hitchin systems
$D_{\bar{A}}=\partial_{\bar{z}}+\bar{A}\,.$ (A.8)
The quotient
$Bun(G)=\{D_{\bar{A}}\}/\mathcal{G}(G)\,,$ (A.9)
where
$\mathcal{G}(G)=C^{\infty}Maps(\Sigma\to G)\,,$ (A.10)
is the moduli space of the holomorphic $G$ bundles over $\Sigma$.
The Higgs field $\Phi$ is a section of the bundle of endomorphisms $E\to E\otimes\mathcal{K}_{\Sigma}$, where $\mathcal{K}_{\Sigma}$ is the canonical class. The $G$-Higgs bundle is the pair
$\mathcal{H}=(D_{\bar{A}}\,,~{}\Phi)\,.$ (A.11)
It is the cotangent bundle $T^{*}\mathcal{A}$ to the affine space of holomorphic structures $\mathcal{A}=\{D_{\bar{A}}\}$ on $E$. Therefore, the Higgs bundle can be endowed with the canonical symplectic form
$\Omega=\int_{\Sigma}|d^{2}z|(\delta\Phi\stackrel{{\scriptstyle\wedge}}{{,}}\delta\bar{A})\,,$ (A.12)
where (,)(~{},~{})( , ) is a fixed invariant form on the Lie algebra ${\mathfrak{g}}^{\mathbb{C}}$, and $\stackrel{{\scriptstyle\wedge}}{{,}}$ means the inner product in the symplectic form.
[…]
$D_{\bar{A}}\Phi=0\,.$ (A.13)
This equation defines the holomorphic structure on the Higgs bundle $\mathcal{H}$ (A.11). Let $\mathcal{C}_{0}$ be the set of solutions of (A.13). The symplectic quotient
$\mathcal{M}^{H}(G)=\mathcal{H}//\mathcal{G}(G)\sim\mathcal{C}_{0}/\mathcal{G}(G)$ (A.14)
is the phase space of the Hitchin integrable systems [12, 13, 17] with the Poisson brackets coming from (A.12).
5.3 Affine Hitchin systems
To pass to 2d integrable systems we consider the affine version of the Higgs bundle. For this purpose we replace the group $G$ on the group $\hat{L}(G)$ (A.4), (A.5).
Let us consider an infinite rank $L(G)$-bundle $E^{aff}$ and a line bundle $\mathcal{L}$ over $\Sigma$. Consider also smooth maps $C^{\infty}(\Omega^{(0,1)}(\Sigma)\to\hat{\mathfrak{g}})=\{(\bar{A}(z,\bar{z},y),\bar{\kappa}(z,\bar{z})\}$ (see (A.1)). The components of this map are connection forms on the bundles $E^{aff}$ and $\mathcal{L}$:
$\nabla_{\bar{A}}=\left(\begin{array}[]{c}\partial_{\bar{z}}+\bar{A}(z,\bar{z},y)\\ \partial_{\bar{z}}+\bar{\kappa}(z,\bar{z})\end{array}\right)\,.$ (A.15)
Affine Higgs connection.
Consider the map $C^{\infty}(\mathcal{K}_{\Sigma}\to\hat{L}^{*}({\mathfrak{g}}))$ (A.2) given by
$\Omega=\frac{1}{\pi}\int_{\Sigma}|d^{2}z|\Bigl{(}\langle\delta\Phi\stackrel{{\scriptstyle\wedge}}{{,}}\delta\bar{A}\rangle+\delta{k}\wedge\delta\bar{\kappa}\Bigr{)}\,.$ (A.16)
Marked points.
Introduce two sets of marked points on $\Sigma$
$\mathcal{D}_{1}=\{z_{a}\,,~{}a=1,\ldots,n\}\,,~{}~{}\mathcal{D}_{2}=\{w_{b}\,,~{}b=1,\ldots,m\}\,.$ (A.17)
${\mathcal{O}}(p_{a}^{(0)},c_{a}^{(0)})=\{S^{a}(p_{a}^{(0)},c_{a}^{(0)})=g^{-1}p_{a}^{(0)}g+c_{a}^{(0)}g^{-1}\partial g\,,~{}g\in L(G)\}\,.$ (A.18)
The orbit (passing through a point $p_{a}^{(0)}$, and with the central charge $c_{a}^{(0)}$) is the coset space ${\mathcal{O}}(p^{(0)},c^{(0)})\sim L(G)/H$ for $c^{(0)}\neq 0$, and ${\mathcal{O}}(p^{(0)},0)\sim L(G)/L(H)$, where $H$ is the Cartan subgroup of $G$. Thereby the orbit can be described by the elements of these quotient:
$S^{a}\to g_{a}\in L(G)/H\,,~{}{\rm or}~{}g_{a}\in L(G)/L(H)\,.$ (A.19)
The invariants defining the orbit ${\mathcal{O}}(p^{(0)},c^{(0)})$ are the conjugacy classes of the monodromy operator $M$ corresponding to the operator $c^{(0)}\partial+S$. Therefore, there is a one-to-one correspondence between the set of $L(G)$-orbits in the coalgebra $\hat{L}^{*}({\mathfrak{g}})$ (A.2) and the set of conjugacy classes in the group $G$.
Symplectic form.
The symplectic form $\omega_{a}$ on the orbit $\mathcal{O}_{a}$ (A.18) takes the form
$\omega_{a}=\frac{1}{2\pi\imath}\oint\delta\langle S^{a}(p_{a}^{(0)},c_{a}^{(0)}),g^{-1}\delta g\rangle\frac{dy}{y}\,.$ (A.20)
The corresponding Poisson brackets are
$\{S^{a}_{\alpha}(y),S^{a}_{\beta}(y^{\prime})\}=\delta(y/{y^{\prime}})\sum\limits_{\gamma}c^{\gamma}_{\alpha\beta}S^{a}_{\gamma}(x)+c^{(0)}\kappa_{\alpha\beta}\partial\delta(y/{y^{\prime}})\,,$ (A.21)
where $\kappa_{\alpha\beta}$ is the defined above invariant form on ${\mathfrak{g}}$, $c^{\gamma}_{\alpha\beta}$ are structure constants in ${\mathfrak{g}}$, and $\alpha,\beta,\gamma$ here are indices in some basis in ${\mathfrak{g}}$. The form $\omega_{a}$ (A.20) is invariant under transformations $g\to gf$, $f\in L(G)$. The corresponding moment map is $S^{a}(p_{a}^{(0)},c_{a}^{(0)})$.
Define the symplectic form $\Omega$ on the space of fields
$\mathcal{H}(G)=(\bar{A}\,,\,\Phi\,,\,{k}\,,\,\bar{\kappa}\,,\,\cup_{a=1}^{n}S^{a})\,.$ (A.22)
Taking into account (A.16) and (A.20) we have
$\Omega=\frac{1}{\pi}\int_{\Sigma}|d^{2}z|\Bigl{(}\langle\delta\Phi\stackrel{{\scriptstyle\wedge}}{{,}}\delta\bar{A}\rangle+\delta{k}\wedge\delta\bar{\kappa}-\sum_{a=1}^{n}\omega_{a}\delta(z-z_{a},\bar{z}-\bar{z}_{a})\Bigr{)}\,,$ (A.23)
Gauge symmetries – symplectomorphisms.
Define the map
$G(\Sigma)=C^{\infty}(\Sigma\to L(G))=\{f(z,\bar{z},y)\}\,.$ (A.24)
The gauge group has two components
$\hat{\cal G}:=(G(\Sigma),\mathcal{G}_{1})\,,~{}~{}\mathcal{G}_{1}=\exp\Big{(}\varepsilon_{1}(z,\bar{z})\Big{)}\,,~{}\varepsilon_{1}(z,\bar{z})\in C^{\infty}(\Sigma)\,,$ (A.25)
as well as its Lie algebra
The coalgebras $L^{*}_{0}$ and $L^{*}_{1}$ are distributions on these spaces.
The actions of $L_{0}$ and $L_{1}$ on the fields take the form
$\begin{array}[]{ll}1.&\delta_{\epsilon}\bar{A}(z,\bar{z},x)=-\partial_{\bar{z}}\epsilon+[\epsilon,\bar{A}]\,,\\ 2.&\delta_{\epsilon}\bar{\kappa}(z,\bar{z})=\langle\bar{A}\partial\epsilon\rangle\,.\end{array}$ (A.26)
These formulas are the infinite-dimensional version of (A.6). Also we have
$\begin{array}[]{ll}1.&\delta_{\varepsilon_{1}}\bar{A}(z,\bar{z},x)=0\,,\\ 2.&\delta_{\varepsilon_{1}}\bar{\kappa}(z,\bar{z})=-\partial_{\bar{z}}\varepsilon_{1}\,.\end{array}$ (A.27)
Taking into account the coadjoint action (A.7) we find
$\delta_{\epsilon}\Phi={k}\partial\epsilon+[\Phi,\epsilon]\,,~{}~{}\delta_{\epsilon}{k}=0\,,$ (A.28)
$\delta_{\varepsilon_{1}}\Phi=0\,,~{}~{}\delta_{\varepsilon_{1}}{k}=0\,.$ (A.29)
$\delta_{\epsilon}g_{a}=g_{a}\epsilon(z_{a},\bar{z}_{a})\,,~{}~{}\delta_{\varepsilon_{1}}g_{a}=0\,.$ (A.30)
Symplectic quotient.
The moment maps $\mu_{j}\,:\,\mathcal{H}(G)\to Lie^{*}(\hat{\cal G}^{G})$, $j=1,2$ are defined by the actions (A.26)-(A.30):
$\mu_{0}=\partial_{\bar{z}}\Phi-{k}\partial\bar{A}+[\bar{A},\Phi]-\sum_{a=1}^{n}S^{a}\delta(z-z_{a},\bar{z}-\bar{z}_{a})\in L_{0}^{*}\,,$ (A.31)
$\mu_{1}=\partial_{\bar{z}}{k}\in L_{1}^{*}=(C^{\infty})^{*}(\Sigma)\,.$ (A.32)
Consider the set of the marked points $\mathcal{D}_{2}$ (A.17). Take
$\mu_{1}=\mu_{1}^{0}\,,~{}~{}\mu_{1}^{0}=\sum_{b=1}^{m}s_{b}\delta(z-w_{b},\bar{z}-\bar{w}_{b})\,,$ (A.33)
where
$\sum_{b=1}^{m}s_{b}=0\,.$ (A.34)
$1.\ \mu_{0}=0\,,\qquad 2.\ \mu_{1}=\mu_{1}^{0}\,.$ (A.35)
In this way we come to the symplectic quotient
This is the set of the gauge equivalent solutions of the moment equations (A.35).